BOOK REVIEW

Wilson, Robin J. Stamping Through Mathematics. Springer–Verlog, Inc., New York, 2001, pp. 126.

An absolutely delightful voyage through the history of mathematics is presented here by Robin Wilson, the well–known British mathematics historian and popularist and author of the “Stamp Corner” section in the Mathematics Intelligencer. Wilson presents his narration with the unique visual aid of postage stamps. These stamps, issued by practically every country (even the Unites States!) in the world, are bedazzling in their color, shape, and design. This is a book that just leaps out at you with its brilliant display of mathematics.

This is by no means meant to be a textbook on mathematics history. It does not contain any exercises, and does not contain adequate text on historical matters. Its purpose, though, is to present a slide show on how the countries of the world have recognized the importance of mathematics by depicting individuals, mathematical achievements and results, and areas of human endeavor that have been influenced by mathematics. We see, for example, stamps that depict such luminaries as Pythagoras, Euclid, Archimedes, Zu Changzhi, al–Biruni, Copernicus, Descartes, Pascal, Newton, Liebniz, Euler, Buffon, Lagrange, Galois, Chebyshev, Kovalevskaya, Bessel, Gauss, Russell, Einstein, Babbage, Poincare, Ramanujan, and von Neumann. Other stamps commemorate famous formulas and identities like a^{2} + b^{2} = c^{2}, E = mc^{2}, λ = h/mv, e^{ln(n)} = n, 1 + 1 = 2, or

f(a) = \frac{1}{2πi} ∫_{Γ} \frac{f(z)}{z − a} dz

while some depict the importance of map making, navigational instruments, astronomy, counting and calculating machines, and applications to physics, art, recreation, and education.

The book leads the reader on a mathematical journey through time. Subdivided into 55 sections, each with one page of text and one page of 6 to 8 stamps, mathematics history is illustrated from the early Greeks and their devotion to geometry, to the Chinese, and then Islamic mathematics. This is followed by 14th and 15th century Europeans and their quest to explore the world, to the explosion of 17th and 18th century mathematics, highlighted by many great French mathematicians, and by algebra, calculus, Newton, Gauss, Euler, and number theory. Twentieth century developments including computer science, painting, nature, quantum theory, and the metric system round out the book.

In addition to a useful bibliography at the end of the book the author lists every single stamp (almost 400) featured in the book by its origin, date of issue, and stamp classification number so philatelists can locate its exact position in the data bank of world wide issuance of stamps. Moderately priced at $25, the book is a delightful addition to one’s library, not only because it is light and fascinating reading but can also be used to spark student interest in any mathematics course.

Reviewed by Thomas P. Dence Department of Mathematics and Computer Science Ashland University Ashland, OH 44839